OpenStudy (anonymous):

Subtract, simplify if possible. 4-v 3v-7 ---- - ----- v-8 8-v

OpenStudy (anonymous):

\[\frac{a}{b}-\frac{c}{d}=\frac{a d-b c}{b d} \]\[\frac{2 v-3}{v-8} \]